OpenStudy (anonymous):

a rocket goes at 49,000,000 inches per hour. how many miles per second??

OpenStudy (anonymous):

square miles?

OpenStudy (anonymous):

\[49 \times10^6 miles/hr \times 1hr/3600\sec \times 1.578\times10^-5/1inch = 0.2148 miles/\sec\]

OpenStudy (stormfire1):

1 mile = 5280ft * 12in/ft = 63360 inches So you have 49000000/63360 = ~773.36 mph Divide that 3600 seconds in an hour and you get 0.2148 miles/sec